# Simulation-based inference using surjective sequential neural likelihood estimation

### v2

@@ Abstract
Neural likelihood estimation methods for simulation-based inference can suffer from performance degradation when the modelled data is very high-dimensional or lies along a lower-dimensional manifold due to the inability of the density estimator to accurately infer a density function. We present Surjective Sequential Neural Likelihood (SSNL) estimation, a novel member in the family of methods for simulation-based inference. SSNL fits a dimensionality-reducing surjective normalizing flow model and uses it as a surrogate likelihood function which allows for computational inference via Markov chain Monte Carlo or variational Bayes methods. Among other benefits, SSNL avoids the requirement to manually craft summary statistics for inference of high-dimensional data sets, since the lower-dimensional representation is computed simultaneously to learning the likelihood and without additional computational overhead. We evaluate SSNL on a wide variety of experiments including two challenging real-world examples from the astrophysics and neuroscience literatures, and show that it generally outperforms state-of-the-art methods.
simulation-based inference, Bayesian inference, normalizing flows, density estimation
@@ 1 Introduction
Probabilistic models are essential tools in the sciences to draw conclusions from data and to make decisions based on computational inferences. In the natural sciences, especially in disciplines such as biology, physics or psychology, Bayesian inference has become particularly popular due to its ability both to quantify uncertainty in parameter values using probability statements, and to incorporate prior knowledge about quantities of interest. Bayesian statistics infers statistical parameters $\theta$ by conditioning the elicited prior distributions $p(\theta)$ on data ${y}$ yielding a posterior distribution $p(\theta|{y})\propto p(y|\theta)p(\theta)$. If the likelihood function $p({y}|\theta)$ is available, i.e., tractable to compute, conventional Bayesian inference using Markov chain Monte Carlo or variational methods can be used for parameter inference (Brooks et al., 2011; Wainwright & Jordan, 2008). However, for many scientific hypotheses, the likelihood is not easy to compute and the experimenter merely has access to a simulator function ${sim}(\theta)$ that can generate synthetic data conditionally on a parameter configuration $\theta$.
@@
In the latter case, an emergent family of methods collectively called simulation-based inference (SBI, Cranmer et al. (2020)) has been proposed. Traditionally, approximate Bayesian computation (ABC, Sisson et al. (2018)), and most successfully sequential Monte Carlo ABC (SMC-ABC; e.g., Beaumont et al. (2009); Lenormand et al. (2013)) or simulated annealing ABC (SABC; e.g., Albert et al. (2015)), has been used to infer approximate posterior distributions (Pritchard et al., 1999; Ratmann et al., 2007). More recently, methods that are based on neural density or density-ratio estimation have found increased application in the natural sciences due to their reduced computational cost and convincing off-the-shelf performance (Brehmer et al., 2018; Delaunoy et al., 2020; Gonçalves et al., 2020; Hermans et al., 2021; Brehmer, 2021; Dax et al., 2021). Among these, several branches of methods exist. Likelihood-based methods (Papamakarios et al., 2019; Glöckler et al., 2022) fit a surrogate model for the likelihood function using neural density estimators (Papamakarios et al., 2021) which allows to do conventional Bayesian inference and which has been shown to bring significant performance advantages in comparison to ABC methods with the same computational budged. Cranmer et al. (2015); Durkan et al. (2020); Hermans et al. (2020); Thomas et al. (2022); Delaunoy et al. (2022) and Miller et al. (2022) developed similar methods that instead target the likelihood-to-evidence ratio rather than the likelihood, while Papamakarios & Murray (2016); Lueckmann et al. (2017); Greenberg et al. (2019) and Deistler et al. (2022) developed methods that try to approximate the posterior distribution directly.
@@
In the case of likelihood-based methods, the accuracy of posterior inferences might suffer due to the inability of neural density estimators to correctly approximate the surrogate likelihoods, e.g., if the data is very high-dimensional, the data are embedded in a low-dimensional manifold but lie in a higher-dimensional ambient space, or when the data contain non-informative data dimensions (Fefferman et al., 2016; Kingma & Dhariwal, 2018; Greenberg et al., 2019; Cunningham et al., 2020; Dai & Seljak, 2021; Klein et al., 2021).
@@
To overcome this limitation, we present a new method for simulation-based inference which we call Surjective Sequential Neural Likelihood (SSNL) estimation. SSNL uses a surjective dimensionality-reducing normalizing flow to model the surrogate likelihood of a Bayesian model by that allowing improved density estimation and consequently improved posterior inference, especially in high-dimensional settings. We evaluate SSNL on several experiments from the SBI, astrophysics and neuroscience literatures and demonstrate that it achieves state-of-the-art performance in comparison to recent methods such as SNL (Papamakarios et al., 2019), automatic posterior transformation (SNPE-C, Greenberg et al. (2019)) or sequential contrastive neural ratio estimation (SNRE-C, Miller et al. (2022)). Conversely, we also demonstrate negative examples when our method should, in theory and empirically, not have a performance gain in comparison to the recent literature.
The structure of the manuscript is as follows. Section 2 describes relevant background for the introduced method and experiments. Section 3 introduces Surjective Sequential Neural Likelihood estimation. Section 4 shows the results on the conducted experiments and demonstrates the performance of the method. In Section 5, related and recent literature is reviewed. Finally, Section 6 summarizes and discusses our contributions and reflects on future research directions in SBI.
@@ 2 Background
Given a parameter configuration $\theta\sim p(\theta)$, a simulator function $sim(\theta)$ is a computer program or experimental procedure that simulates an observable ${y}\leftarrow sim(\theta)$. Apart from stochasticity produced by $p(\theta)$, the simulator might be making use of another source of endogenous randomness. The simulator defines, albeit implicitly, a conditional probability distribution $p({y}|\theta)$ to which the modeller does not have access or which they cannot evaluate in reasonable time. The goal of SBI is to infer the posterior probability distribution $p(\theta|{y})\propto p({y}|\theta)p(\theta)$ for, in many cases, a specific observation ${y}_{0}$ (Cranmer et al., 2020). In the following, we introduce relevant background on density estimation with normalizing flows and neural likelihood methods (background on neural posterior and ratio methods can be found in Appendix A).
@@ 2.1 Density estimation using normalizing flows
Sequential density-based SBI methods, such as Greenberg et al. (2019); Papamakarios et al. (2019); Deistler et al. (2022), use conditional normalizing flows to fit a surrogate model to either approximate the intractable likelihood or posterior. Normalizing flows (NFs, Papamakarios et al. (2021)) model a probability distribution via a pushforward measure as
@@
where $\det J_{k}=\det\frac{\partial f_{k}}{\partial{z}_{k-1}}$ is the determinant of the Jacobian matrix of a forward transformation $f_{k}$ which is typically parameterized with a neural network with weights $\phi_{k}$, and $p\left({z}_{0}\right)$ is some base distribution that has a density that can be evaluated exactly, for instance, a spherical multivariate Gaussian. The forward transformations $f=(f_{1},\dots,f_{K})$ are a sequence of $K$ diffeomorphisms which are applied consecutively to compute $y=z_{K}=f_{K}\circ\dots\circ f_{2}\circ f_{1}(z_{0})$. The two densities $q$ and $p$ are related by the multiplicative term $\det J_{k}$ which is needed to account for the change-of-volume induced by $f_{k}$ and which is termed likelihood contribution in Nielsen et al. (2020) and Klein et al. (2021). The diffeomorphisms $f_{k}$ are required to be dimensionality-preserving and invertible to be able to both evaluate the probability of a data set and to draw samples.
2.2 Sequential neural likelihood estimation
Sequential neural likelihood estimation (SNL, Papamakarios et al. (2019)) iteratively fits a density estimator to approximate the likelihood via $q_{f}({y}|\theta)\approx p({y}|\theta)$. SNL proceeds in $R$ rounds: in the first round, $r=1$, a prior sample $\theta_{n}\sim p(\theta)$ of size $N$ is drawn and then used to simulate data points $y_{n}\leftarrow sim(\theta_{n})$ yielding the data set $\mathcal{D}=\{({y}_{n},\theta_{n})\}^{r}_{1\dots N}$. The simulated data is used to train a conditional normalizing flow by maximizing the expected probability $\mathbb{E}_{\mathcal{D}}\left[q_{f}({y}|\theta)\right]$. Having access to a likelihood approximation, posterior realizations can be generated either by sampling from $\hat{p}^{r}(\theta|{y}_{0})\propto q_{f}({y}_{0}|\theta)p(\theta)$ via Markov chain Monte Carlo or via optimization by fitting a variational approximation to the approximate posterior. SNL then uses the surrogate posterior as proposal prior distribution for the next round, $r+1$, i.e., it draws a new sample of parameters $\theta_{n}\sim\hat{p}^{r}(\theta|{y}_{0})$ which are then used to simulate a new batch of pairs $\{({y}_{n},\theta_{n})\}^{r+1}_{1\dots N}$. The data sets from the previous rounds and the current round are then appended together and a new model is trained on the entire data set. With an infinite computational, or experimental, budget and a sufficiently flexible density estimator $q_{f}$, SNL converges to the desired posterior distribution $p(\theta|{y}_{0})$. In practice, however, the budget is typically limited, the number simulations should be held as small as possible, and finite data leads to inaccurate approximations of likelihood functions.
@@ 3 Surjective Sequential Neural Likelihood Estimation
Surjective Sequential Neural Likelihood (SSNL) estimation approximates the intractable likelihood $p(y|\theta)$ function of a Bayesian model $p(\theta|y)$ while simultaneously embedding the data in a low dimensional manifold using dimensionality-reducing surjections. It then uses this approximation to compute the joint density of the model which allows generating samples from the posterior.
@@
We assume that if the data lie in a high-dimensional ambient space, which is the case for many real-world data sets or time series data, embedding them in a lower-dimensional space should improve likelihood estimation and consequently posterior inference. We parameterize the surrogate likelihood $q_{f}(y|\theta)$ as a combination of $K$ conditional dimensionality-preserving and dimensionality-reducing normalizing flows (for ease of notation, we in the following drop sub- and superscripts wherever it does not decrease clarity).
Specifically, we model the transformations using surjective normalizing flows (Nielsen et al., 2020; Klein et al., 2021) by splitting a data vector ${y}\in\mathbb{R}^{P}$ into two subvectors ${y}=\left[{y}_{-},{y}_{+}\right]^{T}$ and defining the pullback, i.e., from data to embedding space, only on one of these subvectors
${z}=f^{-1}({y}_{+})$
We then define the respective forward transformation generatively as
${y}=\begin{bmatrix}{y}_{+}=f({z})\\ {y}_{-}\sim p({y}_{-}|{z})\end{bmatrix}$
@@
where $y_{+},{z}\in\mathbb{R}^{Q}$ and $Q$ is a hyperparameter. The subvectors are obtained by (arbitrarily) defining two disjoint permutations $\pi_{+}\cup\pi_{-}=\{1,\dots,P\}$, $\pi_{+}\cap\pi_{-}=\emptyset$, and then setting $y_{+}=[y_{\pi_{+}(1)},\dots,y_{\pi_{+}(Q)}]^{T}$ and $y_{-}=[y_{\pi_{-}(1)},\dots,y_{\pi_{-}(P-Q)}]^{T}$. More concretely, we define the inverse mapping as $f^{-1}({y}_{+};{y}_{-},\theta)$ as a conditional normalizing flow (and its forward as $f(z;{y}_{-},\theta)$), and introduce an additional conditional density $p\left({y}_{-}|f^{-1}({y}_{+};{y}_{-},\theta),\theta\right)$ which we parameterize using a neural network.
For a surjective normalizing flow, the "likelihood contribution" becomes the product of the Jacobian determinant of the transformation acting on ${y}_{+}$ and the conditional probability $p\left(y_{-}|f^{-1}({y}_{+};{y}_{-},\theta),\theta\right)$, such that for a surjective flow with only a single transformation the probability of an observation is computed as
$q_{f}({y}|\theta)=p\left({z}\right)p({y}_{-}|{z},\theta)\big{|}\det J\big{|}^{-1}$
@@
where ${z}=f^{-1}({y}_{+};{y}_{-},\theta)$ and $\det J=\det\frac{\partial f(\cdot;{y}_{-},\theta)}{\partial{z}}$ is again the Jacobian determinant of the forward transformation acting on the low-dimensional vector ${z}$ (see Appendix B for a detailed derivation of the surjection layer and appendical Figure 6 for an illustration). Note that this representation strictly extends the one by Klein et al. (2021), since here we additionally require conditioning on the parameter vector $\theta$.
Analogously to multi-layered bijective flows (Equation (1)), the conditional density of a normalizing flow that consists of $K$ dimensionality-reducing layers has the following form:
@@
where $z_{k,-}$ and $z_{k,+}$ are subvectors of $z_{k}$ that have been constructed as above and $J_{k}=\frac{\partial f_{k}(\cdot;{z}_{k,-},\theta)}{\partial{z}_{k-1,+}}$ is the Jacobian of the $k$th transformation $f_{k}$.
SSNL models the likelihood used for simulation-based-inference as a composition of dimensionality-preserving and -reducing layers:
@@
where $\mathcal{K}_{\text{pres}}$ and $\mathcal{K}_{\text{red}}$ represent sets of indexes for dimensionality-preserving and -reducing flow layers, respectively. For instance, for a total of $K=5$ normalizing flow layers, one could alternate between bijections and surjections by setting the sets $\mathcal{K}_{\text{pres}}=\{1,3,5\}$ and $\mathcal{K}_{\text{red}}=\{2,4\}$. We commonly parameterize $f$ using masked autoregressive flows (Germain et al., 2015; Papamakarios et al., 2017), but in practice, e.g., coupling flows (Dinh et al., 2015, 2017) or affine spline flows (Durkan et al., 2019), are also possible.
The dimensionality-reducing flow is fully deterministic in the pullback direction, i.e., in the case of likelihood estimation, but requires sampling from the conditional $p\left({y}_{-}|f^{-1}({y}_{+};{y}_{-},\theta),\theta\right)$ during the forward transformation and, hence, has additional stochastic components other than the base distribution $p({z}_{0})$. For our setting, i.e., density estimation, this is however not a limitation.
The lower-dimensional embedding of SSNL solves previous issues of neural likelihood methods when scaled to high-dimensional data. In addition, through the dimensionality-reduction the flows require less trainable parameters, which empirically speeds up computation such that more of the computational budget can be used for the simulator or more expressive architectures. The embedding acts, albeit only conceptually, as a collection of summary statistics which consequently replaces the need of manually defining them.
Like other sequential methods, SSNL is trained in $R$ rounds where in every round a new proposal posterior is defined. The proposal posterior can be either sampled from using MCMC methods or approximated with another conditional distribution using variational inference (Algorithm 1).
@@
  Inputs: observation ${y}_{0}$, prior distribution $p(\theta)$, surjective normalizing flow $q_{f}({y}|\theta)$, simulations per round $N$, number of rounds $R$
  Outputs: approximate posterior distribution $\hat{p}^{R}(\theta|{y}_{0})$
  Initialize proposal $\hat{p}^{0}(\theta|{y}_{0})\leftarrow p(\theta)$, data set $\mathcal{D}=\{\}$
@@
     for $n\leftarrow 1,\dots,N$ do
        Sample $\theta_{n}\sim\hat{p}^{r-1}(\theta\mid{y}_{0})$
@@
     Set $\hat{p}^{r}(\theta|{y}_{0})\propto q_{f}({y}_{0}|\theta)p(\theta)$

### v3

# Simulation-based Inference for High-dimensional Data using Surjective Sequential Neural Likelihood Estimation

@@ Abstract
Neural likelihood estimation methods for simulation-based inference can suffer from performance degradation when the modeled data is very high-dimensional or lies along a lower-dimensional manifold, which is due to the inability of the density estimator to accurately estimate a density function. We present Surjective Sequential Neural Likelihood (SSNL) estimation, a novel member in the family of methods for simulation-based inference (SBI). SSNL fits a dimensionality-reducing surjective normalizing flow model and uses it as a surrogate likelihood function, which allows for computational inference via Markov chain Monte Carlo or variational Bayes methods. Among other benefits, SSNL avoids the requirement to manually craft summary statistics for inference of high-dimensional data sets, since the lower-dimensional representation is computed simultaneously with learning the likelihood and without additional computational overhead. We evaluate SSNL on a wide variety of experiments, including two challenging real-world examples from the astrophysics and neuroscience literatures, and show that it either outperforms or is on par with state-of-the-art methods, making it an excellent off-the-shelf estimator for SBI for high-dimensional data sets.
@@ 1 Introduction
In the natural sciences, especially in disciplines such as biology and physics, Bayesian inference is becoming increasingly popular due to its ability both to quantify uncertainty in parameter values and to incorporate prior knowledge about quantities of interest. Bayesian statistics infers the posterior distribution $p(\theta|{y})\propto p(y|\theta)p(\theta)$ of statistical parameters $\theta$ by conditioning a prior distributions $p(\theta)$ on data ${y}$. If the likelihood function $p({y}|\theta)$ is available, i.e., tractable to compute, conventional Bayesian inference using Markov chain Monte Carlo or variational methods can be used for parameter inference (Brooks et al.,, 2011; Wainwright and Jordan,, 2008). However, for many scientific hypotheses, the likelihood is not easy to compute and the experimenter merely has access to a simulator function ${sim}(\theta)$ that can generate synthetic data conditionally on a parameter configuration $\theta$.
@@
In the latter case, an emergent family of methods collectively called simulation-based inference (SBI, Cranmer et al., (2020)) has been proposed. Traditionally, approximate Bayesian computation (ABC, Sisson et al., (2018)), and most successfully sequential Monte Carlo ABC (SMC-ABC; e.g., Beaumont et al., (2009); Lenormand et al., (2013)) or simulated annealing ABC (SABC; e.g., Albert et al., (2015)), has been used to infer approximate posterior distributions (Pritchard et al.,, 1999; Ratmann et al.,, 2007). More recently, methods that are based on neural density or density-ratio estimation have found increased application in the natural sciences due to their reduced computational cost and convincing inferential accuracy (Brehmer et al.,, 2018; Delaunoy et al.,, 2020; Gonçalves et al.,, 2020; Hermans et al.,, 2021; Brehmer,, 2021; Dax et al.,, 2021). Among these, several branches of methods exist. Likelihood-based methods (Papamakarios et al.,, 2019; Glöckler et al.,, 2022) fit a surrogate model for the likelihood function using neural density estimators (Papamakarios et al.,, 2021) which allows to do conventional Bayesian inference and which has been shown to bring significant performance advantages in comparison to ABC methods with the same computational budget. Cranmer et al., (2015); Durkan et al., (2020); Hermans et al., (2020); Thomas et al., (2022); Delaunoy et al., (2022); Miller et al., (2022) developed similar methods that instead target the likelihood-to-evidence ratio rather than the likelihood, while Papamakarios and Murray, (2016); Lueckmann et al., (2017); Greenberg et al., (2019); Deistler et al., (2022); Wildberger et al., (2023); Sharrock et al., (2024) developed methods that try to approximate the posterior distribution directly.
@@
In the case of likelihood-based methods, the accuracy of posterior inferences might suffer due to the inability of neural density estimators to correctly approximate the surrogate likelihoods, e.g., if the data are very high-dimensional or the data are embedded in a low-dimensional manifold but lie in a higher-dimensional ambient space (Fefferman et al.,, 2016; Kingma and Dhariwal,, 2018; Greenberg et al.,, 2019; Cunningham et al.,, 2020; Dai and Seljak,, 2021; Klein et al.,, 2021).
@@
To overcome this limitation, we present a new method for simulation-based inference which we call Surjective Sequential Neural Likelihood (SSNL) estimation. SSNL uses a surjective dimensionality-reducing normalizing flow to model the surrogate likelihood of a Bayesian model by that allowing improved density estimation and consequently improved posterior inferences. We evaluate SSNL on multiple experiments from the SBI, astrophysics and neuroscience literatures and demonstrate that it achieves superior performance in comparison to state-of-the-art methods. Conversely, we also demonstrate negative examples when our method should, in theory and empirically, not have a performance gain.
(a) Bijection.
(b) Surjection.
Figure 1: Conditional bijective and surjective flow layers illustrated with masked autoregressive flows. (a) A bijective flow layer $f^{-1}(y;\theta)$ transforms an input $y_{j}$ conditional on all previous values $y_{<j}$ and a parameter vector $\theta$ using a conditioner $c$ and transformer $\tau$. (b) The surjective flow layer $f^{-1}(y_{+};y_{-},\theta)$ first splits the vector into two components $y_{+}$ and $y_{-}$ and then uses the component $y_{-}$ as additional conditioning variable. The implementations for conditioner and transformer remain the same as for the bijection. To evaluate the likelihood, a surjective layer additionally computes the conditional density $p(y_{-}|z,\theta)$ (which computationally is done after the transform).
@@ 2 Background
Given prior parameter values $\theta\sim p(\theta)$, a simulator function $sim(\theta)$ is a computer program or experimental procedure that can simulate an observation ${y}\leftarrow sim(\theta)$. Apart from stochasticity produced by $p(\theta)$, the simulator might be making use of another source of endogenous randomness. The simulator defines, albeit implicitly, a conditional probability distribution $p({y}|\theta)$ to which the modeller does not have access or which they cannot evaluate in reasonable time. The goal of SBI is to infer the posterior distribution $p(\theta|{y})\propto p({y}|\theta)p(\theta)$ using synthetic data $\{(y_{n},\theta_{n})\}_{n=1}^{N}$ generated from the prior model $p(\theta)$ and simulator $sim(\theta)$. Typically, the total simulation budget $N$ is limited and the posterior for a specific observation ${y}_{\text{obs}}$ is the target of inference (Cranmer et al.,, 2020). In the following, we introduce relevant background on density estimation with normalizing flows and neural likelihood methods (background on neural posterior and ratio estimation methods can be found in Appendix A).
@@ 2.1 Density estimation using normalizing flows
Sequential density-based SBI methods (e.g., Greenberg et al., (2019); Papamakarios et al., (2019); Deistler et al., (2022)) use conditional normalizing flows to fit a surrogate model to either approximate the intractable likelihood or posterior. Normalizing flows (NFs, Papamakarios et al., (2021)) model a probability distribution via a pushforward measure as
@@
where $\det J_{k}=\det\frac{\partial f_{k}}{\partial{z}_{k-1}}$ is the determinant of the Jacobian matrix of a forward transformation $f_{k}$ which is typically parameterized with a neural network, and $p\left({z}_{0}\right)$ is some base distribution that has a density that can be evaluated exactly, for instance, a spherical multivariate Gaussian. The forward transformations $f=(f_{1},\dots,f_{K})$ are a sequence of $K$ diffeomorphisms which are applied consecutively to compute $y=z_{K}=f_{K}\circ\dots\circ f_{2}\circ f_{1}(z_{0})$. The two densities $q$ and $p$ are related by the multiplicative terms $\det J_{k}$ which are needed to account for the change-of-volume induced by $f_{k}$ and which are termed likelihood contribution in Nielsen et al., (2020) and Klein et al., (2021). The diffeomorphisms $f_{k}$ are required to be dimensionality-preserving and invertible to be able to both evaluate the probability of a data point and to draw samples. Particularly, in autoregressive flows (Kingma et al.,, 2016; Papamakarios et al.,, 2017; Germain et al.,, 2015) each transformation $f_{k}$ admits a Jacobian determinant which is efficient to compute and which can be decomposed into an autoregressive conditioner $c_{i}$ and an invertible transformer $\tau_{i}$ as
$z_{k,i}=\tau_{i}(z_{k-1,i},c_{i}(z_{k-1,<i}))$
where all $c_{i}$ can be computed jointly using a masked neural network (Figure 1a).
2.2 Neural likelihood estimation
Sequential neural likelihood estimation (SNL, Papamakarios et al., (2019)) iteratively fits a density estimator to approximate the likelihood via $q_{f}({y}|\theta)\approx p({y}|\theta)$. SNL proceeds in $R$ rounds distributing the total simulation budget $N$ evenly in each of these: in the first round, $r=1$, a prior sample $\theta_{n}\sim p(\theta)$ of size $N_{R}=N/R$ is drawn and used to simulate data points $y_{n}\leftarrow sim(\theta_{n})$ yielding the data set $\mathcal{D}=\{({y}_{n},\theta_{n})\}^{r}_{1\dots N_{R}}$. The simulated data is used to train a conditional normalizing flow by maximizing the expected probability $\mathbb{E}_{\mathcal{D}}\left[q_{f}({y}|\theta)\right]$. Having access to a likelihood approximation, posterior realizations can be generated either by sampling from $\hat{p}^{r}(\theta|{y}_{\text{obs}})\propto q_{f}({y}_{\text{obs}}|\theta)p(\theta)$ via Markov chain Monte Carlo or via optimization by fitting a variational approximation to the approximate posterior. SNL then uses the surrogate posterior as proposal prior distribution for the next round, $r+1$, i.e., it draws a new sample of parameters $\theta_{n}\sim\hat{p}^{r}(\theta|{y}_{\text{obs}})$ which are then used to simulate a new batch of pairs $\{({y}_{n},\theta_{n})\}^{r+1}_{1\dots N_{R}}$. The data sets from the previous rounds and the current round are then appended together and a new model is trained on the entire data set. With an infinite simulation budget and a sufficiently flexible density estimator $q_{f}$, SNL converges to the desired posterior distribution $p(\theta|{y}_{\text{obs}})$. The simulation budget is in practice typically limited, the number simulations should be held as small as possible, and finite data leads to inaccurate approximations to the likelihood functions.
@@ 3 Surjective Sequential Neural Likelihood Estimation
Surjective Sequential Neural Likelihood (SSNL) estimation approximates the intractable likelihood $p(y|\theta)$ function of a Bayesian model $p(\theta|y)$ while simultaneously embedding the data in a lower-dimensional space using dimensionality-reducing surjective flows. We assume that if the data lie in a high-dimensional ambient space, which is the case for many real-world data sets like time series data, embedding them in a lower-dimensional space should improve likelihood estimation and consequently posterior inference.
We motivate the derivation of the surjective flow layer using the holistic generative framework of Nielsen et al., (2020) which models the log-probability of a $P$-dimensional data point ${y}$ as
$\log p({y})\simeq\log p\left({z}\right)+V({y},{z})+E({y},{z}),\quad{z}\sim q({z}|{y})$
where $q({z}|{y})$ is some amortized (variational) distribution, ${z}$ is a latent variable with distribution $p({z})$, $V({y},{z})=\log\tfrac{p(y|z)}{q(z|y)}$ is denoted likelihood contribution term and $E({y},{z})=\log\tfrac{q(z|y)}{p(z|y)}$ is a bound looseness term. Intringuingly, for inference surjections, i.e., the kind of flow layers we are considering here, the likelihood contribution can be calculated as
where $p({y}|{z})$ is a conditional density, $h^{-1}:\mathcal{Y}\rightarrow\mathcal{Z}$ is a dimensionality-reducing mapping, and where we for convenience of notation denote with $h:\mathcal{Z}\rightarrow\mathcal{Y}$ a right inverse function to $h^{-1}$ such that $h^{-1}\circ h=Id_{\mathcal{Z}}$. Critically, for surjective normalizing flows, the bound looseness equals $E({y},{z})=0$ if a right inverse function $h$ exists.
@@
We design a conditional surjection layer for dimensionality-reduction as follows (Figure 1b for a graphical overview). We first split the data vector ${y}\in\mathbb{R}^{P}$ into two subvectors ${y}=\left[{y}_{-},{y}_{+}\right]^{T}$ where $y_{+}\in\mathbb{R}^{Q}$ and $Q$ is a hyperparameter. The subvectors are obtained by (arbitrarily) defining two disjoint permutations $\pi_{+}\cup\pi_{-}=\{1,\dots,P\}$, $\pi_{+}\cap\pi_{-}=\emptyset$, and then setting $y_{+}=[y_{\pi_{+}(1)},\dots,y_{\pi_{+}(Q)}]^{T}$ and $y_{-}=[y_{\pi_{-}(1)},\dots,y_{\pi_{-}(P-Q)}]^{T}$. We then construct a conditional normalizing flow $f(z;y_{-},\theta)$ (i.e., conditional on $y_{-}$ and $\theta$) and its inverse $f^{-1}(y_{+};y_{-},\theta)$ and define
$q({z}|{y})$ $=\delta\left({z}-f^{-1}(y_{+};y_{-},\theta)\right)$
where
is the Jacobian of the inverse mapping (see Appendix B for details). Using this result and the conditional distribution $p({y}|{z})=p({y}_{-}|{z},\theta)$ the likelihood contribution for a surjection layer becomes
$=$ $\int\delta\left({z}-f^{-1}({y}_{+};y_{-},\theta)\right)$
$\quad\log\frac{p({y}_{-}|{z},\theta)}{\delta\left({z}-f^{-1}({y}_{+};y_{-},\theta)\right)}\mathrm{d}{z}$
where we used the change of variables $\tilde{y}^{+}=f(z;y_{-},\theta)$ yielding $\mathrm{d}\tilde{y}^{+}=\mathrm{d}z|\det J^{-1}|^{-1}$. The likelihood of an observation using a surjective flow is consequently the product of three terms:
$p\left({z}\right)p({y}_{-}|{z},\theta)\big{|}\det J\big{|}^{-1}$
@@
where $p(z)$ is a base distribution, ${z}=f^{-1}({y}_{+};{y}_{-},\theta)$ and $\det J=\det\frac{\partial f(\cdot;{y}_{-},\theta)}{\partial{z}}$ is again the Jacobian determinant of the forward transformation acting on the lower-dimensional vector ${z}$ (see Appendix B for a detailed derivation of the surjection layer). Note that this representation strictly extends the one by Klein et al., (2021), since here we construct flows that are conditioned on the parameter vector $\theta$. Analogously to multi-layered bijective flows (Equation (1)), the conditional density of a normalizing flow that consists of $K$ dimensionality-reducing layers has the following form:
@@
where $z_{k,-}$ and $z_{k,+}$ are subvectors of $z_{k}$ that have been constructed as above and $J_{k}=\frac{\partial f_{k}(\cdot;{z}_{k,-},\theta)}{\partial{z}_{k-1,+}}$ is the Jacobian of the $k$th surjective transformation $f_{k}$.
For simulation-based-inference, we model the likelihood estimator $q_{f}({y}|\theta)$ as a composition of dimensionality-preserving and -reducing layers:
@@
where $\mathcal{K}_{\text{pres}}$ and $\mathcal{K}_{\text{red}}$ represent sets of indexes for dimensionality-preserving and -reducing flow layers, respectively. For instance, for a total of $K=5$ normalizing flow layers, one could alternate between bijections and surjections by setting the sets $\mathcal{K}_{\text{pres}}=\{1,3,5\}$ and $\mathcal{K}_{\text{red}}=\{2,4\}$. Here, we parameterize $f$ using masked autoregressive flows but in general any flow architecture, such as coupling flows (Dinh et al.,, 2015, 2017), neural spline flows (Durkan et al.,, 2019) or neural autoregressive flows (Huang et al.,, 2018; De Cao et al.,, 2020), is possible.
The dimensionality-reducing flow is fully deterministic in the pullback direction, i.e., in the case of likelihood estimation, but requires sampling from the conditional $p\left({y}_{-}|z,\theta\right)$ during the forward transformation and, hence, has additional stochastic components other than the base distribution $p({z}_{0})$. For our setting, i.e., density estimation, this is however not a limitation.
The lower-dimensional embedding of SSNL solves previous issues of neural likelihood methods when scaled to high-dimensional data sets. In addition, through the dimensionality-reduction the flows require less trainable parameters, which can speed up computation such that more of the computational budget can be used for the simulator or more expressive architectures. The embedding acts, albeit only conceptually, as a collection of summary statistics which consequently replaces the need of manually defining them.
Like other sequential methods, SSNL is trained in $R$ rounds where in every round a new proposal posterior is defined. The proposal posterior can be either sampled from using MCMC methods or approximated with another conditional distribution using variational inference (see Algorithm 1).
@@
  Inputs: observation ${y}_{\text{obs}}$, prior distribution $p(\theta)$, surjective normalizing flow $q_{f}({y}|\theta)$, simulations per round $N_{R}$, number of rounds $R$
  Outputs: approximate posterior distribution $\hat{p}^{R}(\theta|{y}_{\text{obs}})$
  Initialize proposal $\hat{p}^{0}(\theta|{y}_{\text{obs}})\leftarrow p(\theta)$, data set $\mathcal{D}=\{\}$
@@
     for $n\leftarrow 1,\dots,N_{R}$ do
        Sample $\theta_{n}\sim\hat{p}^{r-1}(\theta|{y}_{\text{obs}})$
@@
     Set $\hat{p}^{r}(\theta|{y}_{\text{obs}})\propto q_{f}({y}_{\text{obs}}|\theta)p(\theta)$